# Code Equivalence, Point Set Equivalence, and Polynomial Isomorphism

### v1

@@ 1. Introduction
In this paper we study reductions of the LCE problem to other problems, in particular to the Polynomial Isomorphism (PI) problem. Further polynomial time reductions of this problem considered before were a reduction to the lattice isomorphism problem (see [4]) and to the permutation equivalence problem (see [38]). Notice, however, that the latter reduction is not polynomial time in the sense considered here, because we require polynomial time in $(n,k,\log(q))$, whereas [38] only provides a polynomial time algorithm in $(n,k,q)$.
@@
In order to check the latter property, we introduce our final tool, the Macaulay inverse system. Starting with the divided power algebra ${\mathcal{D}}=\underline{\mathop{\rm Hom}\nolimits}_{{\mathbb{F}}_{q}}(P,{\mathbb{F}}_{q})={\mathbb{F}}_{q}[\pi_{1},\dots,\pi_{k}]$, where $P={\mathbb{F}}_{q}[x_{1},\dots,x_{k}]$ and $\pi_{i}$ is the projection to $x_{i}$, we consider the action $\circ$ of $P$ on $D$ by contraction and define for every homogeneous ideal $J\subseteq P$ its Macaulay inverse system as $J^{\perp}=\{\psi\in{\mathcal{D}}\mid f\circ\psi=0$ for all $f\in J\}$. By a theorem of F.S. Macaulay, the inverse system of an Artinian Gorenstein algebra with socle degree $d$ is a principal ideal generated by a homogeneous polynomial of degree $d$ (see Thm.  7.3) and vice versa. Since this duality is equivariant with respect to the action of $\mathop{\rm GL}\nolimits_{k}({\mathbb{F}}_{q})$ by linear coordinate changes (see Prop. 7.4), we reduce the LCE and PSE problem to the Polynomial Isomorphism (PI) problem for the Macaulay inverse polynomials $\varPhi_{{\mathbb{X}}}$ and $\varPhi_{{\mathbb{X}}^{\prime}}$ of $\mathbb{X}$} and $\mathbb{X}$$^{\prime}$, respectively. The latter problem is to to check whether these two homogeneous polynomials of degree $2r_{\mathbb{X}}-1$ differ only by a homogeneous linear change of coordinates, and if this is the case, to find the matrix defining that linear change of coordinates (see Thm. 8.1).
@@
Another setting in which we have a polynomial time reduction is the case of indecomposable iso-dual codes, or equivalently, of indecomposable self-associated point sets. Namely, we can use the fact that the Artinian reduction $\overline{R}=R/\langle\ell\rangle$ of the homogeneous coordinate ring of $\mathbb{X}$} is an Artinian Gorenstein ring with socle degree 3. Thus the computation of the corresponding Macaulay inverse system is achieved in polynomial time, and we have reduced the LCE search problem to the PI search problem (see Prop. 9.1 and Cor. 9.2).
@@
Unless explicitly noted otherwise, the notation in this paper follows the books [30] and [31]. For basic definitions and facts about set of points in projective spaces, we refer the readers to [31, Sect. 6.3].
@@ 2. Linear Code Equivalence and Point Set Equivalence
The following two lemmas show that the hypothesis that the codes under consideration are projective linear codes is no serious restriction for the study of the LCE problem. (Similar lemmas are shown in [13, Lemmas 3.4 and 3.5].)
@@
The proof is analogous to the proof of [13, Lemma 3.5].
@@
Next we prove (b). As in [13, Lemma 3.4], we see that the number of columns of $G$ which corresponds to a certain point in ${\mathbb{P}}^{k-1}$ equals the number of columns of $G^{\prime}$ with this property. This implies $n=n^{\prime}$, and that the columns which are canceled in $G$ and $G^{\prime}$ correspond to each other one to one. Therefore, if we consider $G^{\prime}=A\cdot G\cdot D\cdot P$ and delete in the matrix DPD\,P the row-column pairs corresponding to the canceled columns, we get $\overline{G}^{\prime}=A\cdot\overline{G}\cdot\overline{D\;P}$, and this yields the claim.
Finally, we show (c). As before, the columns which have been canceled in $G$ and $G^{\prime}$ correspond to each other one to one. By multiplying $G$ and $G^{\prime}$ on the right with suitable permutation matrices and arguing by induction, we may assume that $G$ and $G^{\prime}$ arise from $\overline{G}$ and $\overline{G}^{\prime}$, respectively, by appending a column with is a non-zero multiple of the last column. By multiplying $G$ and $G^{\prime}$ on the right by a suitable invertible diagonal matrix, we may even assume that the last two columns of both $G$ and $G^{\prime}$ are equal. Starting from $\overline{G}^{\prime}=A\cdot\overline{G}\cdot\overline{D}\cdot\overline{P}$, it now sufficies to append the row $(0,\dots,0,1)$ and the column $(0,\dots,0,1)^{\rm tr}$ to $\overline{D}$ and to $\overline{P}$ in order to get matrices D,PD,\,P such that $G^{\prime}=A\cdot G\cdot D\cdot P$, as required. ∎
@@
In other words, we assume that the set $\mathbb{X}$} is not a blocking set in ${\mathbb{P}}^{k-1}\!$. For most of the associated point sets of the codes in this paper, this assumption is satisfied. However, if one wants to study a code whose associated point set does not satisfy it, one can enlarge the base field ${\mathbb{F}}_{q}\subseteq{\mathbb{F}}_{q^{f}}$ with $f\geq 2$ and choose a suitable linear form $\ell$ which is not defined over ${\mathbb{F}}_{q}$. Under this base field extension, the relevant invariants of the code and the point set, such as the parameters of the code, or the Hilbert function of the coordinate ring of the point set, do not change. Moreover, if two codes $C,C^{\prime}\subseteq{\mathbb{F}}_{q}^{n}$ are equivalent as ${\mathbb{F}}_{q^{f}}$-linear codes in $({\mathbb{F}}_{q^{f}})^{n}$, the matrices defining this equivalence are automatically contained in $\mathop{\rm GL}\nolimits_{n}({\mathbb{F}}_{q})$, i.e., the the codes are equivalent over ${\mathbb{F}}_{q}$. Thus it suffices to study the code equivalence problem over the extension field.
@@
The Cayley-Bacharach property can also be formulated using the following notion. For $i\in\{1,\dots,n\}$, a polynomial $f\in P$ is called a separator (or sometimes an indicator function) of $p_{i}$ in $\mathbb{X}$} if $f(p_{i})=1$ and $f(p_{j})=0$ for $j\neq i$ (see also [20] or [31, Sect. 6.3]). Here we use the coordinate tuple of $p_{i}$ with respect to the affine space $D_{+}(L)$ to define $f(p_{i})$. More precisely, we use the coordinate system $(L,y_{1},\dots,y_{k-1})$ of ${\mathbb{P}}^{k-1}$ and the homogeneous coordinates $p_{i}=(1:\bar{p}_{i1}:\cdots\bar{p}_{i\,k{-}1})$ with respect to this coordinate system.
@@
For $i=1,\dots,n$, we have $\mathop{\rm sepdeg}\nolimits_{{\mathbb{X}}}(p_{i})\leq r_{\mathbb{X}}$. The set of points $\mathbb{X}$} has the Cayley-Bacharach property if and only if $\mathop{\rm sepdeg}\nolimits_{{\mathbb{X}}}(p_{i})=r_{\mathbb{X}}$ for $i=1,\dots,n$ (see [20] or [31, Tutorial 88]).
@@ 3. The Canonical Module
In the above setting, a presentation of the canonical module $\omega_{R}$ can be calculated as follows (see [32, Sect. 4.5]). From the ${\mathbb{F}}_{q}[\ell]$-basis $\mathcal{O}_{\sigma}(\bar{I}_{\mathbb{X}})=\{t_{1},\dots,t_{n}\}$ of $R$ we calculate, in polynomial time, the multiplication matrices of $x_{1},\dots,x_{k}$ in this basis. Their transposed matrices are the multiplication matrices of the canonical module $\omega_{R}={\mathbb{F}}_{q}[\ell]\,t_{1}^{\ast}\oplus\cdots\oplus{\mathbb{F}}_{q}[\ell]\,t_{n}^{\ast}$.
Next we use these matrices to derive a presentation $\omega_{R}\cong\bigoplus_{i=1}^{n}R\,e_{i}/U$ where $U$ is generated by elements of the form $x_{i}e_{j}-\sum_{\nu=1}^{n}c_{\nu}e_{\nu}$ with $c_{\nu}\in{\mathbb{F}}_{q}[\ell]$. Lastly, we minimize this presentation using the graded version of Nakayama’s Lemma (see [30, Prop. 1.7.15]).
@@
The equivalence of (b) and (c) was shown in [19, Prop. 5.5]. Notice that the set of points $\mathbb{X}$} is non-degenerate, since $\mathop{\rm rk}\nolimits(G)=k$. ∎
@@ 4. Iso-Dual Codes and Self-Associated Sets of Points
Our goal in this section is to reinterpret the dual code of $C$ and the concept of iso-dual codes in terms of the associated sets of points. Let $G$ be a generator matrix of $C$ and ${\mathbb{X}}=\{p_{1},\dots,p_{n}\}$ the associated set of ${\mathbb{F}}_{q}$-rational points in ${\mathbb{P}}^{k-1}$ defined by the columns of $G$. The associated sets of points of $C$ and $C^{\perp}$ are related using the following construction (see [19, Def. 1.1]).
@@
First we show that (a) is equivalent to (b). Since $C$ is iso-dual, we know that $n=2k$. As the rank of $G$ is $k$, the points of $\mathbb{X}$} span ${\mathbb{P}}^{k-1}$. This implies $\mathop{\rm HF}\nolimits_{\mathbb{X}}(1)=k+1$. By Proposition 3.4, the condition in (a) that $C$ is indecomposable is equivalent to the condition that $\omega_{R}$ is generated in degrees $\leq-1$. Hence we can apply [19, Thm. 7.2]. It yields that (a) is equivalent to the condition that $\mathbb{X}$} is arithmetically Gorenstein. In particular, this forces the Hilbert function of $\mathbb{X}$} to be symmetric, and thus we get the claimed function $\Delta\mathop{\rm HF}\nolimits_{\mathbb{X}}$.
The equivalence of (b) and (c) is well-known in commutative algebra (see [23, Satz 6.14] or [22, Prop. 2.1.3]). In our setting, it also follows by combining Prop. 2.12 and [20, Thm. 3.5]. The stated Hilbert function of $\omega_{R}$ follows from the above formula. ∎
@@ 5. The Canonical Ideal
Using the identification $\omega_{R}=\underline{\mathop{\rm Hom}\nolimits}_{{\mathbb{F}}_{q}[\ell]}(R,{\mathbb{F}}_{q}[\ell])(-1)$, the canonical module of $R$ can now be embedded as an ideal of $R$, as the next result shows (see [28, Prop. 1.9]).
@@
The canonical ideal of $\mathbb{X}$} depends on the choice of $\ell\in R_{1}$, of course. Moreover, the canonical module can also be embedded into higher degree components of $R$. The next proposition describes these non-uniquenesses explicitly. A similar result was shown in [9, Thm. 3.4].
@@
Claim (a) was shown in [28, Prop. 1.8]. Claim (b) is an immediate consequence of the description of the homogeneous components and the multiplication of $R$ in degrees $\geq r_{\mathbb{X}}$ given above. ∎
@@
First we prove (a). The set $\{t_{1},\dots,t_{n}\}$ is an ${\mathbb{F}}_{q}[\ell]$-basis of $R$ by [31, Thm. 4.3.22]. To write $f$ in terms of this basis, it suffices to compute the normal form of $f$ with respect to $I_{\mathbb{X}}$ and the extension $\overline{\sigma}$ of $\sigma$ (see [31, Sect. 4.3]). Then $\{t_{1}^{\ast},\dots,t_{n}^{\ast}\}$ is an ${\mathbb{F}}_{q}[\ell]$-basis of $\omega_{R}=\underline{\mathop{\rm Hom}\nolimits}_{{\mathbb{F}}_{q}[\ell]}(R,{\mathbb{F}}_{q}[\ell])(-1)$ and the claim follows from $t_{i}^{\ast}(t_{j})=\delta_{ij}$ for $i,j=1,\dots,n$.
@@
Since $\ell\in R_{1}$ is a non-zerodivisor for $R$, the Artinian reduction $\overline{R}=R/\langle\ell\rangle$ of $R$ satisfies $\mathop{\rm HF}\nolimits_{\overline{R}}(i)=\mathop{\rm HF}\nolimits_{R}(i)-\mathop{\rm HF}\nolimits_{R}(i-1)\neq 0$ if and only if $i\in\{0,\dots,r_{\mathbb{X}}\}$. The value $\Delta=\mathop{\rm HF}\nolimits_{\overline{R}}(r_{\mathbb{X}})$ is called the last difference of $\mathop{\rm HF}\nolimits_{R}$. Then the residue classes $\bar{f}_{1},\dots,\bar{f}_{n}$ of the separators $f_{i}\in R_{r_{\mathbb{X}}}$ generate $\overline{R}_{r_{\mathbb{X}}}$, because $\{f_{1},\dots,f_{n}\}$ is an ${\mathbb{F}}_{q}$-basis of $R_{r_{\mathbb{X}}}$ (see also [20, Prop. 2.13]).
@@
for some elements $g_{j}\in R_{r_{\mathbb{X}}-1}$, and the set $\{g_{1},\dots,g_{n-\Delta}\}$ is an ${\mathbb{F}}_{q}$-basis of $R_{r_{\mathbb{X}}-1}$ (see also [28, Lemma 1.2.b]).
@@
By Theorem 4.4, the set of points $\mathbb{X}$} is arithmetically Gorenstein and $\omega_{R}$ is a graded free $R$-module of rank 1. Moreover, we have $r_{\mathbb{X}}=3$. An element ${\varphi}\in(\omega_{R})_{-2}$ is an $R$-basis of $\omega_{R}$ if and only if ${\varphi}(f_{i})\neq 0$ for $i=1,\dots,n$ (see [20, Lemma 3.4.c]). Then the corresponding element $\pi={\varphi}(f_{1})f_{1}+\cdots+{\varphi}(f_{n})f_{n}$ of $({\mathfrak{J}}_{R/{\mathbb{F}}_{q}[\ell]})_{3}$ is an $R$-basis of the graded free $R$-module ${\mathfrak{J}}_{R/{\mathbb{F}}_{q}[\ell]}$ of rank 1. As $\mathbb{X}$} has the Cayley-Bacharach property, the degree of each point is 3, and thus $f_{i}\notin\ell\,R_{2}$ for $i=1,\dots,n$. Here we have $\dim_{{\mathbb{F}}_{q}}(\overline{R}_{3})=\Delta\mathop{\rm HF}\nolimits_{\mathbb{X}}(3)=1$ and we can write $\bar{f_{i}}=\beta_{i}\bar{f}_{1}$ with $\beta_{i}\in{\mathbb{F}}_{q}\setminus\{0\}$ for $i=2,\dots,n$. Thus, if we let ${\varphi}(f_{1})=1$ and ${\varphi}(f_{i})=\beta_{i}$ for $i=2,\dots,n$, then ${\varphi}(f_{i}-\beta_{i}f_{1})=0$ implies ${\varphi}(R_{2})=0$ and $\varphi$} is an $R$-basis of $\omega_{R}$. Consequently, the element $\pi=f_{1}+\beta_{2}f_{2}+\cdots+\beta_{n}f_{n}$ is an $R$-basis of ${\mathfrak{J}}_{R/{\mathbb{F}}_{q}[\ell]}$. ∎

### v2

@@ 1. Introduction
In this paper we study reductions of the LCE problem to other problems, in particular to the Polynomial Isomorphism (PI) problem. Further polynomial time reductions of this problem considered before were a reduction to the lattice isomorphism problem (see [6]) and to the permutation equivalence problem (see [46]). Notice, however, that the latter reduction is not polynomial time in the sense considered here, because we require polynomial time in $(n,k,\log(q))$, whereas [46] only provides a polynomial time algorithm in $(n,k,q)$. Moreover, in [29, Prop. 3.3] the authors construct a reduction of the LCE problem to the 3-Tensor Isomorphism (3-TI) problem which has been shown to reduce to the PI problem in degree 3 over fields satisfying certain assumptions on their characteristic (see [2], [3] and [28]). Note that these reductions result in the study of cubic polynomials in significantly more intdeterminates than the ones considered here. A somewhat better reduction from 3-TI to 3-PI follows from [30], but it still yields cubic polynomials involving significantly more indeterminates than necessary in this paper.
@@
In order to check the latter property, we introduce our final tool, the Macaulay inverse system. Starting with the divided power algebra ${\mathcal{D}}=\underline{\mathop{\rm Hom}\nolimits}_{{\mathbb{F}}_{q}}(P,{\mathbb{F}}_{q})={\mathbb{F}}_{q}[\pi_{1},\dots,\pi_{k}]$, where $P={\mathbb{F}}_{q}[x_{1},\dots,x_{k}]$ and $\pi_{i}$ is the projection to $x_{i}$, we consider the action $\circ$ of $P$ on $D$ by contraction and define for every homogeneous ideal $J\subseteq P$ its Macaulay inverse system as $J^{\perp}=\{\psi\in{\mathcal{D}}\mid f\circ\psi=0$ for all $f\in J\}$. By a theorem of F.S. Macaulay, the inverse system of an Artinian Gorenstein algebra with socle degree $d$ is a principal ideal generated by a homogeneous polynomial of degree $d$ (see Thm.  7.3) and vice versa. Since this duality is equivariant with respect to the action of $\mathop{\rm GL}\nolimits_{k}({\mathbb{F}}_{q})$ by linear coordinate changes (see Prop. 7.4), we reduce the LCE and PSE problem to the Polynomial Isomorphism (PI) problem for the Macaulay inverse polynomials $\varPhi_{{\mathbb{X}}}$ and $\varPhi_{{\mathbb{X}}^{\prime}}$ of $\mathbb{X}$} and $\mathbb{X}$$^{\prime}$, respectively. The latter problem is to check whether these two homogeneous polynomials of degree $2r_{\mathbb{X}}-1$ differ only by a homogeneous linear change of coordinates, and if this is the case, to find the matrix defining that linear change of coordinates (see Thm. 8.1).
@@
Another setting in which we have a polynomial time reduction is the case of indecomposable iso-dual codes, or equivalently, of indecomposable self-associated point sets. Namely, we can use the fact that the Artinian reduction $\overline{R}=R/\langle\ell\rangle$ of the homogeneous coordinate ring of $\mathbb{X}$} is an Artinian Gorenstein ring with socle degree 3. Thus the computation of the corresponding Macaulay inverse system is achieved in polynomial time, and we have reduced the LCE search problem to the PI search problem (see Prop. 9.1 and Cor. 9.3).
@@
Unless explicitly noted otherwise, the notation in this paper follows the books [38] and [39]. For basic definitions and facts about set of points in projective spaces, we refer the readers to [39, Sect. 6.3].
@@ 2. Linear Code Equivalence and Point Set Equivalence
The following two lemmas show that the hypothesis that the codes under consideration are projective linear codes is no serious restriction for the study of the LCE problem. (Similar lemmas are shown in [16, Lemmas 3.4 and 3.5].)
@@
The proof is analogous to the proof of [16, Lemma 3.5].
@@
Next we prove (b). As in [16, Lemma 3.4], we see that the number of columns of $G$ which corresponds to a certain point in ${\mathbb{P}}^{k-1}$ equals the number of columns of $G^{\prime}$ with this property. This implies $n=n^{\prime}$, and that the columns which are canceled in $G$ and $G^{\prime}$ correspond to each other one to one. Therefore, if we consider $G^{\prime}=A\cdot G\cdot D\cdot P$ and delete in the matrix DPD\,P the row-column pairs corresponding to the canceled columns, we get $\overline{G}^{\prime}=A\cdot\overline{G}\cdot\overline{D\;P}$, and this yields the claim.
Finally, we show (c). As before, the columns which have been canceled in $G$ and $G^{\prime}$ correspond to each other one to one. By multiplying $G$ and $G^{\prime}$ on the right with suitable permutation matrices and arguing by induction, we may assume that $G$ and $G^{\prime}$ arise from $\overline{G}$ and $\overline{G}^{\prime}$, respectively, by appending a column which is a non-zero multiple of the last column. By multiplying $G$ and $G^{\prime}$ on the right by a suitable invertible diagonal matrix, we may even assume that the last two columns of both $G$ and $G^{\prime}$ are equal. Starting from $\overline{G}^{\prime}=A\cdot\overline{G}\cdot\overline{D}\cdot\overline{P}$, it now sufficies to append the row $(0,\dots,0,1)$ and the column $(0,\dots,0,1)^{\rm tr}$ to $\overline{D}$ and to $\overline{P}$ in order to get matrices D,PD,\,P such that $G^{\prime}=A\cdot G\cdot D\cdot P$, as required. ∎
@@
In other words, we assume that the set $\mathbb{X}$} is not a blocking set in ${\mathbb{P}}^{k-1}\!$. By [9, Thm. 2], a blocking set consists of at least $\frac{q^{k-1}-1}{q-1}$ points. For most of the associated point sets of the codes in this paper, the above assumption is satisfied. However, if one wants to study a code whose associated point set does not satisfy it, one can enlarge the base field ${\mathbb{F}}_{q}\subseteq{\mathbb{F}}_{q^{f}}$ with $f\geq 2$ and choose a suitable linear form $\ell$ which is not defined over ${\mathbb{F}}_{q}$. Under this base field extension, the relevant invariants of the code and the point set, such as the parameters of the code, or the Hilbert function of the coordinate ring of the point set, do not change. Moreover, if two codes $C,C^{\prime}\subseteq{\mathbb{F}}_{q}^{n}$ are equivalent as ${\mathbb{F}}_{q^{f}}$-linear codes in $({\mathbb{F}}_{q^{f}})^{n}$, the matrices defining this equivalence are automatically contained in $\mathop{\rm GL}\nolimits_{n}({\mathbb{F}}_{q})$, i.e., the codes are equivalent over ${\mathbb{F}}_{q}$. Thus it suffices to study the code equivalence problem over the extension field.
@@
The Cayley-Bacharach property can also be formulated using the following notion. For $i\in\{1,\dots,n\}$, a polynomial $f\in P$ is called a separator (or sometimes an indicator function) of $p_{i}$ in $\mathbb{X}$} if $f(p_{i})=1$ and $f(p_{j})=0$ for $j\neq i$ (see also [25] or [39, Sect. 6.3]). Here we use the coordinate tuple of $p_{i}$ with respect to the affine space $D_{+}(L)$ to define $f(p_{i})$. More precisely, we use the coordinate system $(L,y_{1},\dots,y_{k-1})$ of ${\mathbb{P}}^{k-1}$ and the homogeneous coordinates $p_{i}=(1:\bar{p}_{i1}:\cdots\bar{p}_{i\,k{-}1})$ with respect to this coordinate system.
@@
For $i=1,\dots,n$, we have $\mathop{\rm sepdeg}\nolimits_{{\mathbb{X}}}(p_{i})\leq r_{\mathbb{X}}$. The set of points $\mathbb{X}$} has the Cayley-Bacharach property if and only if $\mathop{\rm sepdeg}\nolimits_{{\mathbb{X}}}(p_{i})=r_{\mathbb{X}}$ for $i=1,\dots,n$ (see [25] or [39, Tutorial 88]).
@@ 3. The Canonical Module
In the above setting, a presentation of the canonical module $\omega_{R}$ can be calculated as follows (see [40, Sect. 4.5]). From the ${\mathbb{F}}_{q}[\ell]$-basis $\mathcal{O}_{\sigma}(\bar{I}_{\mathbb{X}})=\{t_{1},\dots,t_{n}\}$ of $R$ we calculate, in polynomial time, the multiplication matrices of $x_{1},\dots,x_{k}$ in this basis. Their transposed matrices are the multiplication matrices of the canonical module $\omega_{R}={\mathbb{F}}_{q}[\ell]\,t_{1}^{\ast}\oplus\cdots\oplus{\mathbb{F}}_{q}[\ell]\,t_{n}^{\ast}$.
Next we use these matrices to derive a presentation $\omega_{R}\cong\bigoplus_{i=1}^{n}R\,e_{i}/U$ where $U$ is generated by elements of the form $x_{i}e_{j}-\sum_{\nu=1}^{n}c_{\nu}e_{\nu}$ with $c_{\nu}\in{\mathbb{F}}_{q}[\ell]$. Lastly, we minimize this presentation using the graded version of Nakayama’s Lemma (see [38, Prop. 1.7.15]).
@@
The equivalence of (b) and (c) was shown in [22, Prop. 5.5]. Notice that the set of points $\mathbb{X}$} is non-degenerate, since $\mathop{\rm rk}\nolimits(G)=k$. ∎
@@ 4. Iso-Dual Codes and Self-Associated Sets of Points
Our goal in this section is to reinterpret the dual code of $C$ and the concept of iso-dual codes in terms of the associated sets of points. Let $G$ be a generator matrix of $C$ and ${\mathbb{X}}=\{p_{1},\dots,p_{n}\}$ the associated set of ${\mathbb{F}}_{q}$-rational points in ${\mathbb{P}}^{k-1}$ defined by the columns of $G$. The associated sets of points of $C$ and $C^{\perp}$ are related using the following construction (see [22, Def. 1.1]).
@@
First we show that (a) is equivalent to (b). Since $C$ is iso-dual, we know that $n=2k$. As the rank of $G$ is $k$, the points of $\mathbb{X}$} span ${\mathbb{P}}^{k-1}$. This implies $\mathop{\rm HF}\nolimits_{\mathbb{X}}(1)=k+1$. By Proposition 3.4, the condition in (a) that $C$ is indecomposable is equivalent to the condition that $\omega_{R}$ is generated in degrees $\leq-1$. Hence we can apply [22, Thm. 7.2]. It yields that (a) is equivalent to the condition that $\mathbb{X}$} is arithmetically Gorenstein. In particular, this forces the Hilbert function of $\mathbb{X}$} to be symmetric, and thus we get the claimed function $\Delta\mathop{\rm HF}\nolimits_{\mathbb{X}}$.
The equivalence of (b) and (c) is well-known in commutative algebra (see [31, Satz 6.14] or [27, Prop. 2.1.3]). In our setting, it also follows by combining Prop. 2.12 and [25, Thm. 3.5]. The stated Hilbert function of $\omega_{R}$ follows from the above formula. ∎
@@ 5. The Canonical Ideal
Using the identification $\omega_{R}=\underline{\mathop{\rm Hom}\nolimits}_{{\mathbb{F}}_{q}[\ell]}(R,{\mathbb{F}}_{q}[\ell])(-1)$, the canonical module of $R$ can now be embedded as an ideal of $R$, as the next result shows (see [36, Prop. 1.9]).
@@
The canonical ideal of $\mathbb{X}$} depends on the choice of $\ell\in R_{1}$, of course. Moreover, the canonical module can also be embedded into higher degree components of $R$. The next proposition describes these non-uniquenesses explicitly. A similar result was shown in [12, Thm. 3.4].
@@
Claim (a) was shown in [36, Prop. 1.8]. Claim (b) is an immediate consequence of the description of the homogeneous components and the multiplication of $R$ in degrees $\geq r_{\mathbb{X}}$ given above. ∎
@@
First we prove (a). The set $\{t_{1},\dots,t_{n}\}$ is an ${\mathbb{F}}_{q}[\ell]$-basis of $R$ by [39, Thm. 4.3.22]. To write $f$ in terms of this basis, it suffices to compute the normal form of $f$ with respect to $I_{\mathbb{X}}$ and the extension $\overline{\sigma}$ of $\sigma$ (see [39, Sect. 4.3]). Then $\{t_{1}^{\ast},\dots,t_{n}^{\ast}\}$ is an ${\mathbb{F}}_{q}[\ell]$-basis of $\omega_{R}=\underline{\mathop{\rm Hom}\nolimits}_{{\mathbb{F}}_{q}[\ell]}(R,{\mathbb{F}}_{q}[\ell])(-1)$ and the claim follows from $t_{i}^{\ast}(t_{j})=\delta_{ij}$ for $i,j=1,\dots,n$.
@@
Since $\ell\in R_{1}$ is a non-zerodivisor for $R$, the Artinian reduction $\overline{R}=R/\langle\ell\rangle$ of $R$ satisfies $\mathop{\rm HF}\nolimits_{\overline{R}}(i)=\mathop{\rm HF}\nolimits_{R}(i)-\mathop{\rm HF}\nolimits_{R}(i-1)\neq 0$ if and only if $i\in\{0,\dots,r_{\mathbb{X}}\}$. The value $\Delta=\mathop{\rm HF}\nolimits_{\overline{R}}(r_{\mathbb{X}})$ is called the last difference of $\mathop{\rm HF}\nolimits_{R}$. Then the residue classes $\bar{f}_{1},\dots,\bar{f}_{n}$ of the separators $f_{i}\in R_{r_{\mathbb{X}}}$ generate $\overline{R}_{r_{\mathbb{X}}}$, because $\{f_{1},\dots,f_{n}\}$ is an ${\mathbb{F}}_{q}$-basis of $R_{r_{\mathbb{X}}}$ (see also [25, Prop. 2.13]).
@@
for some elements $g_{j}\in R_{r_{\mathbb{X}}-1}$, and the set $\{g_{1},\dots,g_{n-\Delta}\}$ is an ${\mathbb{F}}_{q}$-basis of $R_{r_{\mathbb{X}}-1}$ (see also [36, Lemma 1.2.b]).
@@
By Theorem 4.4, the set of points $\mathbb{X}$} is arithmetically Gorenstein and $\omega_{R}$ is a graded free $R$-module of rank 1. Moreover, we have $r_{\mathbb{X}}=3$. An element ${\varphi}\in(\omega_{R})_{-2}$ is an $R$-basis of $\omega_{R}$ if and only if ${\varphi}(f_{i})\neq 0$ for $i=1,\dots,n$ (see [25, Lemma 3.4.c]). Then the corresponding element $\pi={\varphi}(f_{1})f_{1}+\cdots+{\varphi}(f_{n})f_{n}$ of $({\mathfrak{J}}_{R/{\mathbb{F}}_{q}[\ell]})_{3}$ is an $R$-basis of the graded free $R$-module ${\mathfrak{J}}_{R/{\mathbb{F}}_{q}[\ell]}$ of rank 1. As $\mathbb{X}$} has the Cayley-Bacharach property, the degree of each point is 3, and thus $f_{i}\notin\ell\,R_{2}$ for $i=1,\dots,n$. Here we have $\dim_{{\mathbb{F}}_{q}}(\overline{R}_{3})=\Delta\mathop{\rm HF}\nolimits_{\mathbb{X}}(3)=1$ and we can write $\bar{f_{i}}=\beta_{i}\bar{f}_{1}$ with $\beta_{i}\in{\mathbb{F}}_{q}\setminus\{0\}$ for $i=2,\dots,n$. Thus, if we let ${\varphi}(f_{1})=1$ and ${\varphi}(f_{i})=\beta_{i}$ for $i=2,\dots,n$, then ${\varphi}(f_{i}-\beta_{i}f_{1})=0$ implies ${\varphi}(R_{2})=0$ and $\varphi$} is an $R$-basis of $\omega_{R}$. Consequently, the element $\pi=f_{1}+\beta_{2}f_{2}+\cdots+\beta_{n}f_{n}$ is an $R$-basis of ${\mathfrak{J}}_{R/{\mathbb{F}}_{q}[\ell]}$. ∎